OpenStudy (anonymous):

Have a STATE test on Tuesday! Can someone help me study if I have them the sample test? Please? I really need it? (:

OpenStudy (aravindg):

it asks for username and password :O

OpenStudy (anonymous):

I knoq @AravindG I am so sorry! But I do know what I need to study and learn but I suck at math and need help learing it, step by step! If you would like to help me just message me or something? Thanks! :)

OpenStudy (anonymous):

I am just trying my best to meet the requirments of this test.....

OpenStudy (aravindg):

well why should I message you about helping ? I am ready to help already ..Post your questions/doubts I can surely guide you with good inputs

OpenStudy (anonymous):

Okay! Thank you so much! to start off I am a lot behind in math because My school also has sucky percentages of their state test scores so I'm moving schools by the end of this year. and last year (8th grade) I failed the reading, writing, math, and science state tests! So I'm just trying to help myself to at least meet the requirements for this years!

OpenStudy (anonymous):

I started getting behind in Systems of Equations and Inequalities.

OpenStudy (anonymous):

Solving Systems by Graphing Solve each system by graphing. (Is there another way, with an equation, instad of graphing?) y= 2x y= -2x +8 and then I didn't get the one with the fraction either: y= 1/2x + 7 y= 3/2 + 3

OpenStudy (anonymous):

By the way thank you so much! :) @AravindG

OpenStudy (aravindg):

to solve : y= 2x y= -2x +8 you have different methods including graphing method

OpenStudy (aravindg):

easiest method would be to add the two given equations y= 2x + y= -2x +8 ___________ 2y=0+8 y=4

OpenStudy (aravindg):

to get corresponding x value plug the y value in any of the equation above ..For simplicity I am taking first equation y=2x 4=2x x=2

OpenStudy (aravindg):

did you understand that ?

OpenStudy (anonymous):

No.... I don't understand how you got the 4?

OpenStudy (aravindg):

I just added the 2 equations

OpenStudy (anonymous):

Nevermind you adde3d the 2's

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

Sorry! :/ my bad.

OpenStudy (aravindg):

np ..I can understand .Everyone will face difficulty when you start with this topic believe me !

OpenStudy (aravindg):

Its just take practice :)

OpenStudy (anonymous):

haha yes. practice makes perfect. :) But wouldn't that be 4x? and you can't add a neg. to a pos, without having to subtract it.

OpenStudy (aravindg):

ya I just added them 2x+(-2x)=2x-2x=0

OpenStudy (anonymous):

Oh okay!

OpenStudy (aravindg):

by the way khan's academy has good lectures on this topic https://www.khanacademy.org/math/algebra/systems-of-eq-and-ineq

OpenStudy (anonymous):

I've tried that over the summer but I would need a coach if I had any questions so it didnt really work for me because i had a lot of questions

OpenStudy (aravindg):

well then I found this tutorial really simple and illustrative ... http://www.mathwarehouse.com/algebra/linear_equation/systems-of-equation/

OpenStudy (aravindg):

I think you can figure out most of it .When you get doubts just post them here .I can help then

OpenStudy (aravindg):

I highly recommend that tutorial ^^

OpenStudy (anonymous):

Okay .... thanks....

OpenStudy (aravindg):

yw :) infact they have interactive quizzes ..simple explanation ..video at bottom of page ..graphical explanation and much more

OpenStudy (aravindg):

so wishing you great learning experience ..do keep me updated in your progress !

OpenStudy (anonymous):

:) I might be back of confusion so be ready if I do come back! haha. (;

OpenStudy (aravindg):

of course :)